Tohban Report 2016-10-26
Tohban Reports | |
---|---|
Start Date: | 19 October 2016 |
End Date: | 26 October 2016 |
Tohban: | Mitsuo Oka |
Tohban email: | moka@ssl.berkeley.edu |
Next Tohban: | Pascal |
List all reports |
Solar Activity
The solar activity was extremely low, remained at B-level throughout the week. There is only one AR on disk, and so the solar activity is expected to remain low next week.
How many GOES flares occurred?
Flares above B, C, M, X class were 3 0 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 1 0 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 1 0 0 0
There were RHESSI flares/GOES flares 7 / 3 over the time range 18-Oct-16 25-Oct-16
Memory Management
The SSR was at 0% at the end of BGS passes.
Spacecraft Status
The spacecraft has remained stable. Detectors 3 and 9 continued recording events. The cold tips temperature have come down further, and they are currently 117.6 and 115.9 K as of 2016-10-26. The cold plates, 1 and 2 are 142.0 and 140.8 K.
Data Gaps
There were some relatively large data gaps (as shown below) in addition to numerous short (300 s) gaps (not shown).
GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 2016-10-21T19:25:00.000 -- 2016-10-21T19:55:00.000 1800.0000 2016-10-21T20:25:00.000 -- 2016-10-21T20:35:00.000 600.00000 2016-10-21T21:00:00.000 -- 2016-10-21T21:15:00.000 900.00000 2016-10-25T22:40:00.000 -- 2016-10-26T00:00:00.000 4800.0000
GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4 2016-10-23T00:50:00.000 -- 2016-10-23T14:50:00.000 50400.000 2016-10-24T23:30:00.000 -- 2016-10-25T00:00:00.000 1800.0000 2016-10-25T22:40:00.000 -- 2016-10-26T00:00:00.000 4800.0000
GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 2016-10-21T01:40:00.000 -- 2016-10-21T01:55:00.000 900.00000 2016-10-21T10:45:00.000 -- 2016-10-21T10:55:00.000 600.00000 2016-10-21T11:55:00.000 -- 2016-10-21T12:15:00.000 1200.0000 2016-10-21T12:30:00.000 -- 2016-10-21T12:40:00.000 600.00000 2016-10-21T13:00:00.000 -- 2016-10-21T13:20:00.000 1200.0000 2016-10-21T17:30:00.000 -- 2016-10-21T17:40:00.000 600.00000 2016-10-25T02:00:00.000 -- 2016-10-25T02:15:00.000 900.00000 2016-10-25T04:25:00.000 -- 2016-10-26T00:00:00.000 70500.000
Detector issues
Detectors 3 and 9 are powered on, and are recording events. The D9 rear fast threshold was raised from 0x80 to 0xA0.
2016-293-20:04:23 /IDPUTABLE9 OFFSET=REARFASTDAC 2016-293-20:04:33 /IDPULOAD VALUE=0x90
2016-293-20:06:16 /IDPUTABLE9 OFFSET=REARFASTDAC 2016-293-20:06:24 /IDPULOAD VALUE=0xA0
Notes
Spacecraft Management
Decimation | Normal/Vigorous |
HLAT Decimation | No decimation |
Night time data (fronts) | full-time |
Night time data (rears) | full-time |
Require extra passes? | No |
Requirement for moving pointer? | No |
Attenuator operation | Working as expected |
Detector problems? | See notes above. |